OpenStudy (anonymous):

The Sears Tower is 1,454 feet tall. Let the height of the antenna on the tower be represented by the variable a. The total height of the building plus the antenna is 1,707 feet. Which equation represents this problem?

OpenStudy (anonymous):

1454+a=1707

OpenStudy (lilg132):

1707 = 1454 + a

OpenStudy (lilg132):

1707 - a = 1454

OpenStudy (lilg132):

no medal given